# Violations of the null convergence condition in kinematical transitions between singular and regular black holes, horizonless compact objects, and bounces

### v2

@@ IV Discussion
Regarding the Penrose singularity theorem, on the other hand, in an evolving time-dependent situation, an inner horizon does not necessarily need to be a Cauchy horizon, and so it could be compatible with the assumption of global hyperbolicity. In such a situation, it seems one would be forced to violate the NCC in order to reach a singularity-free end point.
@@
We have therefore developed a number of kinematic models that interpolate between “standard physics” (standard black holes or horizonless objects) and regular black holes (assuming spherical symmetry), and have explicitly verified that during the purely kinematic evolution process (no particular dynamics is assumed) violations of the NCC are quite common and apparently unavoidable, although a formal proof of the universality of this behaviour would certainly be desirable.
@@
We feel that this investigation serves to highlight more than ever the pressing need for improved singularity theorems better-adapted to dealing with evolving inner and outer trapping horizons Hawking:2014tga ; Visser:2014zqa ; Ashtekar:2004cn , as well as an improved understanding of existent singularity theorems and their global and geometric convergence conditions. In particular, our investigations emphasize the need for an improved understanding of general conditions under which the formation of timelike singularities — shielded by dynamical inner horizons — is unavoidable. In conclusion, one might say that also in this situation, as has recently happened for mass inflation Carballo-Rubio:2024dca , regular black holes are forcing us to carefully reconsider the dynamical aspects of trapped regions and their horizons. As such, independently of their phenomenological interest, they ought to be considered crucial objects for current and future research in black hole physics (see e.g. Carballo-Rubio:2023mvr ; DiFilippo:2024spj ; Buoninfante:2024oxl ; Visser:2024zkx ; Carballo-Rubio:2025fnc ).

### v3

@@ I Introduction
Although our analysis is purely classical, our results have important implications for the quantum theory associated with the formation of regular black holes and bouncing spacetimes. The necessary NCC violation observed during the transition from a singular state to a regular state can be expected, in any physically realistic scenario, to have an intrinsically quantum origin. While computing the renormalized stess-energy tensor (RSET) in such dynamical situations remains a challenging and unsolved task, our analysis reveals necessary properties required of the RSET in order to induce the formation of a regular final state.
This article is structured as follows. In Section II we explicitly derive the conditions imposed by the NCC in a general dynamical spherically symmetric spacetime. Furthermore we distinguish between the two physically most relevant choices of radial functions suitable for describing either black holes in Subsection II.1, or bouncing spacetimes in Subsection II.2. In Section III we analyze violations of the NCC in kinematical transition models between two stationary geometries. To that end, in Subsection III.1 we first introduce smooth interpolating functions to model such transitions. Subsequently, we analyze violations of the NCC in transitions from singular to regular black holes in Subsection III.2, from regular black holes to horizonless compact objects in Subsection III.3, from singular black holes to bouncing or wormhole spacetimes in Subsection III.4, and from bouncing to wormhole spacetimes in Subsection III.5. We finish with a discussion in Section IV.
@@ IV Discussion
Regarding the Penrose singularity theorem, on the other hand, in an evolving time-dependent situation, an inner horizon does not necessarily need to be a Cauchy horizon, and thereby could be compatible with the assumption of global hyperbolicity. In such a situation, it seems that one would be forced to violate the NCC in order to reach a singularity-free endpoint.
@@
On these grounds we have developed several kinematic models that interpolate between “standard physics” (standard black holes or horizonless objects) and regular black holes (assuming spherical symmetry). and have explicitly verified that during the purely kinematic evolution process (with no particular dynamics is assumed) violations of the NCC are common and seemingly unavoidable, although a formal proof of the universality of this behaviour would certainly be desirable.
Let us emphasize the broad implications and placement of our classical analysis within the context of the formation of regular black holes and bouncing spacetimes in dynamical situations, and more generally within the context of the quantum theory of such regular objects. The NCC violation required for the transition from a singular initial state to a regular final state in a physically realistic scenario can expected to be induced by quantum effects. A complete description and understanding of this NCC violation would require knowledge of the renormalized stress-energy tensor in dynamical situations.
One may consider the viewpoint that NCC violation is secondary to the regularization of black holes, and that it is instead the violation of the Generalized Second Law (GSL) which plays a central role for the quantum theory of regular black holes Wall:2010jtc . Such quantum singularity theorems are however equally inapplicable as the classical singularity theorems for describing the transition to a regular stationary regular black-hole geometry, for which the assumption of global hyperbolicity is violated. One may further object that in physically realistic dynamical situations no Cauchy horizon is present which would violate global hyperbolicity. However, in such situations the applicability of quantum singularity theorems which make use of the GSL is not obvious, due to lack of suitable notions of entropy and theremodynamics quantities in an out-of-equilibriium context.
Let us return to the question of the Cauchy horizon in dynamical situations, and under the impact of quantum effects. It is well-known that Cauchy horizons are subject to a classical instability Carballo-Rubio:2018pmi ; Carballo-Rubio:2021bpr ; DiFilippo:2022qkl , which notably extends to a classical instability of dynamical inner horizons Carballo-Rubio:2024dca .
In addition to the classical instability of Cauchy horizons, there is also a quantum instability arising from the incompatibility between the Unruh state and the presence of stationary inner horizons Hollands:2019whz , see also McMaken:2024tpc ; McMaken:2024fvq ; Balbinot:2023vcm . This illustrates that once a regular black hole has formed, its inner horizon will be subject to an evolving instability. At the present stage the outcome of this evolution is still unknown and moreover it is unclear how such a semiclassical instability could be extended to dynamical inner horizons. Knowledge of the endpoint of such instabilities would in principle allow us to apply our method to this particular final state and determine the renormalized stress-energy tensor (RSET) required to evolve from a singular solution to this endpoint configuration. In the absence of this information we have limited our investigations to the formation of a regular black hole, demonstrating that the interpolation between singular and non-singular solutions necessarily entails a transient violation of the NCC.
@@
Finally let us come back to the classical aspects of our analysis. We feel that this investigation serves to highlight more than ever the pressing need for improved singularity theorems better-adapted to dealing with evolving inner and outer trapping horizons Hawking:2014tga ; Visser:2014zqa ; Ashtekar:2004cn , as well as an improved understanding of existent singularity theorems and their global and geometric convergence conditions. In particular, our investigations emphasize the need for an improved understanding of general conditions under which the formation of timelike singularities — shielded by dynamical inner horizons — is unavoidable. In conclusion, one might say that also in this situation, as has recently happened for mass inflation Carballo-Rubio:2024dca , regular black holes are forcing us to carefully reconsider the dynamical aspects of trapped regions and their horizons. As such, independently of their phenomenological interest, they ought to be considered crucial objects for current and future research in black hole physics (see e.g. Carballo-Rubio:2023mvr ; DiFilippo:2024spj ; Buoninfante:2024oxl ; Visser:2024zkx ; Carballo-Rubio:2025fnc ).